OpenStudy (anonymous):

f(x) = 3x^2 - x - 2 what are the x intercepts ?

sam (.sam.):

x-intercept when y=0 3x^2 - x - 2=0 factor (3x+2)(x-1)=0

sam (.sam.):

x=-2/3 , x=1

OpenStudy (anonymous):

x- intercepts means, what are the values of x when y= 0 so can you solve this: 3x^2 - x - 2=0 so you need to factor like how same did.

OpenStudy (anonymous):

sam is it not x = -1/3

sam (.sam.):

its not x=-1/3 but its x=-2/3 , x=1

OpenStudy (anonymous):

ahh ok, i see